# Sparse Training for Federated Learning with Regularized Error Correction

### v1

@@ IV Experiments
Figure 2: The test accuracy performance is compared for FC and CNN models on MNIST with $E=1$ and $B=\infty$. $1$-FLARE is implemented with a sparsity setting of $R=0.001\%$. FFL and Error Correction are included in the comparison. The uncompressed FedAvg is presented as a benchmark for performance. Remarkably, $1$-FLARE outperforms FFL even when FFL is configured with a sparsity of $R=0.01\%$.
@@
Figure 3: The test accuracy performance is compared for CNN model on digit MNIST with $R=0.001\%$, $p=1,2,4$, $\tau=0.05$, $c=1.1$, 10 clients with $B=\infty$ and $E=4,8,16,32$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction and FFL methods. At the bottom of each figure, $E$ and $log_{10}(R)$ values are plotted to illustrate the sparsity and computation levels of $p$-FLARE compared to FFL. It is evident that $p$-FLARE significantly outperforms Error Correction and FFL in all cases.
@@
Figure 4: The test accuracy performance is compared for FC model on digit MNIST with $R=0.001\%$, $p=1,2,4$, $\tau=0.5$, $c=1.05$, 10 clients with $B=\infty$ and $E=4,8,16,32$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction and FFL methods. At the bottom of each figure, $E$ and $log_{10}(R)$ values are plotted to illustrate the sparsity and computation levels of $p$-FLARE compared to FFL. It is evident that $p$-FLARE outperforms Error Correction and FFL in all cases in terms of convergence rate, demonstrating strong performance in accuracy as well.
Figure 5: Algorithm comparison for FC and CNN models. Each bar indicate when each method reaches Top-1 accuracy of $80\%$. Bars are marked with $X$ and transparent if they did not reach $80\%$ at all. FLARE is implemented with $R=0.001\%$. The superiority of FLARE compared to Error Correction and FFL is evident, even when the sparsity level of FLARE is more aggressive.
@@
Figure 6: The test accuracy performance is compared for VGG models on CIFAR10 with $R=0.001\%$, $p=5$, $\tau=0.05$, $c=1.01$, 4 clients with 12500 examples and $B=500$ and $E=1$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction and FFL methods. It is evident that $p$-FLARE significantly outperforms Error Correction and FFL in all cases.
@@
We conducted simulations for image classification tasks employing five distinct models across different FL settings: FC, CNN, VGG 11, VGG 16, and VGG 19 models. Our initial evaluation focused on assessing the performance of FLARE using the FC and CNN models, trained on the MNIST digit dataset [37]. We extensively tested its performance under various settings to explore its robustness for both balanced and imbalanced data distributions. Subsequently, we scaled up to larger models, including VGG11, VGG16 (138.4M parameters), and VGG19 (143.7M parameters) [38], trained on the CIFAR10 dataset [39]. In all experiments, the models were initialized with the same global model, consistent across experiments, and evaluated on the test set every 10 rounds.
@@
Throughout all experiments, FLARE was configured with a sparsity setting of $R=0.001\%$. We compared our method with uncompressed FedAvg, the classic Error Correction method, and the state-of-the-art FFL algorithm [34]. FFL is an adaptive method that adjusts two variables, $E$ and $R$, to efficiently conduct compression schemes such as Error Correction and further minimize the learning error. It has demonstrated superiority over other recent methods [40], including Qsparse, ATOMO, or AdaComm [41, 42, 43]. Therefore, FFL was selected for performance comparison among existing advanced sparsification methods as it achieves the up-to-date state-of-the-art performance. Since FFL adjusts $E$ and $R$ during training, we include plots of $log_{10}(R)$ and $E$ (if it changes) at the bottom of each figure to verify that our method maintains greater sparsity throughout all rounds.
IV-A Experiment 1: FC and CNN Models on the MNIST Dataset
@@
The FC model consists of three hidden layers, each with 4069 neurons (identical setting as in[33]). The CNN consists of two $5\times 5$ convolution layers, the first with 32 channels, the second with 64, each followed with $2\times 2$ max pooling (identical setting as in[2]). For the FC and CNN experiments, the FL setting consists of 10 clients each holds 600 examples, global test set of 10000, and we use SGD with full batch size ($B$ is extremely large) for consistency. In order to assess the robustness of $p$-FLARE (where $p$ denotes the specific parameter used in FLARE), we conduct tests for $E=1,4,8,16,32$ separately, each with $p=1,2,4$. The results are presented in the same figure, as seen in Figs. 2, 3, 4. The $p$-FLARE for the FC model is set with $\tau=0.5$ and decay constant $c=1.05$, and the CNN experiments are set with $\tau=0.05$ and $c=1.1$. Throughout all experiments, we set $t$ to be the median, implying that pulling is applied to only $50\%$ of all model weights, considered to be the most stale based on their accumulator value.
@@
Fig 2 shows $1$-FLARE for CNN and FC models for the $E=1$ case. The FC model with $1$-FLARE shows to acheive sparsity of $99.999\%$ ($R=0.001\%$) converging with only a negligible delay without any Top-1 accuracy damage. The CNN model with $1$-FLARE shows to follow the curve of the uncompressed FedAvg without any delay, having small accuracy damage. $1$-FLARE outperforms FFL and Error Correction, even when we set FFL with $R=0.01\%$ for both models. On the CNN model, FedAvg achieves Top-1 test accuracy of $0.78$, FLARE achieves $0.73$ while FFL with $R=0.01\%$ achieves $0.67$, meaning our method with one order sparser level achieves greater accuracy performances. On the FC model, FLARE accuracy curve merges with the FedAvg after 400 rounds while FFL with $R=0.01\%$ did not reach this point after 1000 rounds.
@@
Fig 3 and Fig 4 show CNN and FC models results for $E=4,8,16,32$, where we use $p=1,2,4$. In Fig 3, it is evident that CNN training is significantly impacted by sparsification with $R=0.001\%$ using Error Correction and FFL, while $p$-FLARE enables training with a reasonable delay. For example, on $E=32$, $4$-FLARE achieves $0.92$ Top-1 test accuracy, while Error Correction achieves only $0.58$.
In Fig 5, bars are employed to highlight the first round at which Top-1 accuracy reaches $80\%$ for each case. Bars are marked with an $X$ and are transparent when this threshold is not achieved at all. The superiority of FLARE compared to Error Correction and FFL is evident, even when the sparsity level of FLARE is more aggressive.
IV-B Experiment 2: VGG 11, 16, 19 Models on the CIFAR10 Dataset
@@
In this experiment, we validate $p$-FLARE on VGG 11, 16, and 19 models trained on the CIFAR-10 dataset with a sparsity setting of $R=0.001\%$. CIFAR-10 comprises 10 classes of $32\times 32$ images with three RGB channels, partitioned into 4 clients. In all three experiments, each client is equipped with 12,500 examples from the dataset, and we set a global test set of 10,000 examples. We use $E=1$, $B=500$, and employ $p=5$, $\tau=0.05$, and $c=1.01$. Each VGG model is set with a learning rate of $0.002$. We compare the performance of $5$-FLARE with FFL and Error Correction. As shown in Fig.6, $5$-FLARE demonstrates faster convergence and sparsity compared to Error Correction and FFL, with superior overall performances. On VGG16, after 1000 rounds, Error Correction and FFL achieve a test accuracy of $0.3$, while $5$-FLARE merges with FedAvg and reaches a test accuracy of 0.5 after 410 rounds. On VGG 11, FLARE achieves a test accuracy of $0.5$ after 310 rounds, while FFL and Error Correction reach this accuracy level only after 960 and 950 rounds, respectively. After 1000 rounds, FLARE, FFL, and Error Correction achieve test accuracies of 0.65, 0.52, and 0.52, respectively. On VGG19, there is a significant delay for all three methods, with FLARE starting to converge at round 1300, while the other two methods start at 1000 but converge much slower from this point. Specifically, FLARE reaches a test accuracy of 0.5 after 1600 rounds, while FFL achieves this accuracy level only after 2200 rounds. At 3000 rounds, the test accuracy achieved by FLARE, FFL, and Error Correction is 0.67, 0.62, and 0.61, respectively. These experiments serve as a demonstration of FLARE on large-scale DNN models, showcasing its ability to successfully accelerate the sparse training process on complex ML tasks such as VGG models.
IV-C Experiment 3: Imbalanced MNIST Dataset with the FC Model
@@
In this experiment, we assess the performance of $p$-FLARE in imbalanced scenarios of class distributions using the FC model. We partition the MNIST dataset into four distinct configurations: Distributing 1200 digit examples to 5 clients, with each client exclusively receiving examples corresponding to 2, 3, 4, or 5 different digit labels, while maintaining the test set unchanged. This setup introduces varying degrees of class imbalance, with each client possessing only 2 different digits (referred to as imbalance level 2), and dataset labels among clients do not overlap. The FC model is tested for $E=1$ and $R=0.001\%$. The Top-1 test accuracy results are presented in Fig.7. For the 2-level imbalance, after 1000 rounds, FedAvg achieves a test accuracy of 0.83, while $1$-FLARE achieves 0.74, and FFL and Error Correction achieve 0.67. For the 3-level imbalance, FedAvg achieves a test accuracy of 0.82, while $1$-FLARE achieves 0.7, and FFL and Error Correction achieve 0.54. In the case of 4-level imbalance, FedAvg achieves a test accuracy of 0.84, while $1$-FLARE achieves 0.74, and FFL and Error Correction achieve 0.62. Finally, for the 5-level imbalance, FedAvg achieves a test accuracy of 0.83, while $1$-FLARE achieves 0.73, and FFL and Error Correction achieve 0.59. Fig.8 compares all cases based on their Top-1 accuracy achieved by each method after 1000 rounds. These results once again highlight the superiority of FLARE over FFL and Error Correction, demonstrating its robustness across different data distributions.
@@
Figure 8: The test accuracy performance is compared for FC model on MNIST after 1000 rounds at different levels of imbalance in data distributions. FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction and FFL methods. It is evident that $1$-FLARE significantly outperforms Error Correction and FFL in all cases.
IV-D The Contamination Effect
As discussed in Section III, the use of a sufficiently large $p$ may lead to the pulling of irrelevant and unnecessary weights, potentially contaminating updates and accumulators. In this experiment, we explore this contamination effect, illustrating it through the same FC experiment as in previous subsections. This time, we allow each client to minimize $\tilde{f^{i}_{1}}(x,\bar{w};\bar{A}^{i}_{1},\bar{w_{1}})$ at all of its in-round optimization steps in each round (i.e., $p=\infty$). We set no decay ($c=1$) and compare it with Error Correction (each client uses naive $f(x,\bar{w})$), with $R=0.01\%,0.001\%,0.0001\%$, and $E=1,5,10$. The results are presented in Fig. 9. Initially, all cases seem to benefit from the pulling, accelerating training. However, a sudden accuracy drop occurs. Results further suggest that with larger $E$, accuracy drops happen sooner, reinforcing the insight that accumulators are contaminated by the pulling. The faster accuracy drops occur because updates were already compensated for being stale, making the pulling more irrelevant as $E$ grows and contaminating the accumulators. Additionally, it is notable that as $R$ grows larger, contamination occurs sooner as well. This is a second indication that accumulators are affected by the pulling. As training becomes less sparse, updates occur more frequently, leading to earlier contamination. It is important to note that the cases with $E=1$ were not harmed at all, and pulling accelerated convergence at all sparsification levels.
As discussed earlier, the decay of $\tau$ can address the contamination issue, albeit at the cost of reduced training acceleration through pulling. Fig. 10 illustrates the impact of $p=\infty$ and decay with values of $c=1.5$ and $c=1.05$. As observed, decay resolves the accuracy drop, but pulling becomes less effective. With $c=1.5$, training is only slightly accelerated but remains extremely stable. On the other hand, $c=1.05$ provides better pulling benefits for $E=1$ than $c=1.5$ (blue dashed vs. blue solid), but $E=10$ and $E=5$ are adversely affected. By utilizing pulling for only $p$ steps, it becomes feasible to reduce the decay constant, making $p$-FLARE robust and efficient.
Figure 9: The test accuracy performance for FC model on MNIST used to demonstrate the contamination effect by comparing Error Correction with loss $f$, and Error Correction with loss $\tilde{f}$.
Figure 10: The test accuracy performance for FC model on MNIST used to demonstrate that decay mitigates the contamination effect but with cost of performance.
V Conclusion
Error Correction has proven to be a promising technique for alleviating communication overhead in FL systems. However, when sparsity is taken to the extreme, it struggles with staled updates, posing a challenge to maintaining model performance. In Error Correction, clients only exchange the most significant changes, retaining the rest locally for accumulation and releasing them for transmission once they accumulate sufficiently.

### v2

@@ V Experiments
Figure 3: The test accuracy performance is compared for FC and CNN models on MNIST with $E=1$ and $B=\infty$. $1$-FLARE is implemented with a sparsity setting of $R=0.001\%$. FFL, EF21, FedProx and Error Correction are included in the comparison. The uncompressed FedAvg is presented as a benchmark for performance. Remarkably, $1$-FLARE outperforms all other methods even when FFL is configured with a sparsity of $R=0.01\%$.
@@
Figure 4: The test accuracy performance is compared for CNN model on digit MNIST with $R=0.001\%$, $p=1,2,4$, $\tau=0.05$, $c=1.1$, 10 clients with $B=\infty$ and $E=4,8,16,32$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction, FFL, EF21 and FedProx methods. At the bottom of each figure, $E$ and $log_{10}(R)$ values are plotted to illustrate the sparsity and computation levels of $p$-FLARE compared to FFL. It is evident that $p$-FLARE significantly outperforms all other methods in all cases.
@@
Figure 5: The test accuracy performance is compared for FC model on digit MNIST with $R=0.001\%$, $p=1,2,4$, $\tau=0.5$, $c=1.05$, 10 clients with $B=\infty$ and $E=4,8,16,32$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction, FFL, EF21 and FedProx methods. At the bottom of each figure, $E$ and $log_{10}(R)$ values are plotted to illustrate the sparsity and computation levels of $p$-FLARE compared to FFL. It is evident that $p$-FLARE outperforms all other methods in all cases in terms of convergence rate, demonstrating strong performance in accuracy as well.
@@
Figure 6: The test accuracy performance is compared for VGG models on CIFAR10 with $R=0.001\%$, $p=5$, $\tau=0.05$, $c=1.01$, 4 clients with 12500 examples and $B=500$ and $E=1$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction, EF21, FedProx, and FFL methods. It is evident that $p$-FLARE significantly outperforms all other methods in all cases.
@@
We conducted simulations for image classification and text generation tasks employing six distinct models across different FL settings: FC, CNN, VGG 11, VGG 16, VGG 19, and GRU models. Our initial evaluation focused on assessing the performance of FLARE using the FC and CNN models, trained on the MNIST digit dataset [39]. We extensively tested its performance under various settings to explore its robustness for both balanced and imbalanced data distributions. Subsequently, we scaled up to larger models, including VGG11, VGG16 (138.4M parameters), and VGG19 (143.7M parameters) [40] trained on the CIFAR10 dataset [41], and conducted a text generation task based on ”The Complete Works of William Shakespeare” [42]. In all experiments, the models were initialized with the same global model, consistent across experiments, and evaluated on the test set every 10 rounds.
@@
Throughout the experiments, FLARE was configured with a sparsity setting up to $R=0.001\%$. We also conducted additional experiments to evaluate the performance with less aggressive sparsification settings. We compared our method with uncompressed FedAvg, the classic Error Correction method, and the state-of-the-art FFL, EF21 and FedProx algorithms [34, 43, 44]. (i) FFL is an adaptive method that adjusts two variables, $E$ and $R$, to efficiently conduct compression schemes such as Error Correction and further minimize the learning error. It has demonstrated superiority over other recent methods [45], including Qsparse, ATOMO, or AdaComm [46, 47, 48]. Since FFL adjusts $E$ and $R$ during training, we include plots of $log_{10}(R)$ and $E$ (if it changes) at the bottom of each figure to verify that our method maintains greater sparsity throughout all rounds. (ii) EF21  [43] is a newly enhanced version of the standard Error Feedback mechanism. It uses a recursive compressor to reduce compression errors throughout the training process. EF21 has demonstrated superior performance over other methods across a wide range of experiments, highlighting its effectiveness in leveraging error feedback to reduce communication costs. (iii) FedProx  [44] is an FL algorithm that employs dedicated regularization during local training, and therefore a good benchmark for comparison. For our experiments, we integrated Top-K sparsification and error feedback with the regularization techniques of FedProx, ensuring that both FLARE and FedProx share the same objective.
V-A Experiment 1: FC and CNN Models on the MNIST Dataset
@@
The FC model consists of three hidden layers, each with 4069 neurons (identical setting as in[33]). The CNN consists of two $5\times 5$ convolution layers, the first with 32 channels, the second with 64, each followed with $2\times 2$ max pooling (identical setting as in[2]). For the FC and CNN experiments, the FL setting consists of 10 clients each holds 600 examples, global test set of 10000, and we use SGD with full batch size ($B$ is extremely large) for consistency. In order to assess the robustness of $p$-FLARE (where $p$ denotes the specific parameter used in FLARE), we conduct tests for $E=1,4,8,16,32$ separately, each with $p=1,2,4$. The results are presented in the same figure, as seen in Figs. 3, 4, 5. The $p$-FLARE for the FC model is set with $\tau=0.5$ and decay constant $c=1.05$, and the CNN experiments are set with $\tau=0.05$ and $c=1.1$. Throughout all experiments, we set $a_{0}$ to be the median of the accumulated error at absolute values, implying that pulling is applied to only $50\%$ of all model weights, considered to be the most staled based on their accumulator value.
@@
Fig 3 shows $1$-FLARE for CNN and FC models for the $E=1$ case. The FC model with $1$-FLARE shows to achieve sparsity of $99.999\%$ ($R=0.001\%$) converging with only a negligible delay without any Top-1 accuracy damage. The CNN model with $1$-FLARE shows to follow the curve of the uncompressed FedAvg without any delay, having small accuracy damage. $1$-FLARE outperforms FFL, EF21, FedProx and Error Correction, even when we set FFL with $R=0.01\%$ for both models. On the CNN model, FedAvg achieves Top-1 test accuracy of $0.78$, FLARE achieves $0.73$ while FFL with $R=0.01\%$ achieves $0.67$, meaning our method with one order sparser level achieves greater accuracy performances. On the FC model, FLARE accuracy curve merges with the FedAvg after 400 rounds while FFL with $R=0.01\%$ did not reach this point after 1000 rounds.
@@
Fig 4 and Fig 5 show CNN and FC models results for $E=4,8,16,32$, where we use $p=1,2,4$. In Fig 4, it is evident that CNN training is significantly impacted by sparsification with $R=0.001\%$ using any of the compared methods, while $p$-FLARE enables training with a reasonable delay. For example, on $E=32$, $4$-FLARE achieves $0.92$ Top-1 test accuracy, while Error Correction achieves only $0.58$.
V-B Experiment 2: VGG 11, 16, 19 Models on the CIFAR10 Dataset
@@
In this experiment, we validate $p$-FLARE on VGG 11, 16, and 19 models trained on the CIFAR-10 dataset with a sparsity setting of $R=0.001\%$. CIFAR-10 comprises 10 classes of $32\times 32$ images with three RGB channels, partitioned into 4 clients. In all three experiments, each client is equipped with 12,500 examples from the dataset, and we set a global test set of 10,000 examples. We use $E=1$, $B=500$, and employ $p=5$, $\tau=0.05$, and $c=1.01$. Each VGG model is set with a learning rate of $0.002$. We compare the performance of $5$-FLARE with FFL, EF21, FedProx Error Correction. As shown in Fig.6, $5$-FLARE demonstrates faster convergence and sparsity compared to all compared methods, with superior overall performances. On VGG16, after 1000 rounds, all other methods fail to achieve a test accuracy over $0.3$, while $5$-FLARE merges with FedAvg and reaches a test accuracy of 0.5 after 410 rounds. On VGG11, FLARE achieves a test accuracy of $0.5$ after 310 rounds, while all compared methods reach this accuracy level not sooner than 950 rounds. After 1000 rounds, FLARE, FFL, and Error Correction achieve test accuracies of 0.65 and 0.52, respectively. On VGG19, there is a significant delay for all the sparsification methods, with FLARE starting to converge at round 1300, while the others methods start around 1000 but converge much slower from this point. Specifically, FLARE reaches a test accuracy of 0.5 after 1600 rounds, while all others achieve this accuracy level only around 2200 rounds. At 3000 rounds, the test accuracy achieved by FLARE and Error Correction is 0.67 and 0.61, respectively. These experiments serve as a demonstration of FLARE on large-scale DNN models, showcasing its ability to successfully accelerate the sparse training process on complex ML tasks such as VGG models.
V-C Experiment 3: Text generation based on ”The Complete Works of William Shakespeare”
We have conducted text generation experiments based on ”The Complete Works of William Shakespeare” [42], introduced in [2] to simulate a real-world FL task. In this setting, each client represents a different character from Shakespeare’s plays, thus providing unique local text data. This setup creates a realistic FL environment with unbalanced, non-IID data. The model is trained to predict the next character in the text sequence, following the setup in [2], with the following parameters. The FL system consists of $10$ clients, and utilizes a large GRU architecture with $4,022,850$ parameters. Our configuration aligns with the settings described in [49] for $10$ clients, though our model was trained from scratch with a learning rate of $0.05$. For evaluation, we used the same test set proposed by [42]. In these experiments, each client operated with a batch size of $8$, $E=1$ and $R=0.001\%$. FLARE was tested with $\tau=0.05$, $c=1.01$ and $p=\infty$ (i.e. regularization is used every optimization step). As in all other experiments, we set the masking coefficient $a_{0}$ to the median of the accumulated absolute error values. Fig. 8 displays the loss curves on the test set for each method evaluated. Notably, FLARE demonstrates a significant performance advantage over the other algorithms, following FedAvg without any performance degradation.
V-D Experiment 4: Client unavailability
Next, we evaluate how client unavailability affects the convergence and performance of our method. Specifically, we have conducted the experiments with the FC and CNN architectures where, out of 10 total clients, only 7, 5, or 3 randomly selected clients were available during each communication round. For each scenario, we tested with $E=1$. The results, displayed in Figure 7, show that while performance degradation occurs with fewer available clients across all algorithms, as expected, the proposed FLARE algorithm continues to show superior performance compared to the other methods.
(a) FC with 7 available clients
(b) FC with 5 available clients
(c) FC with 3 available clients
(d) CNN with 7 available clients
(e) CNN with 5 available clients
(f) CNN with 3 available clients
Figure 7: The test accuracy performance for client unavailability compared for FC and CNN models on digit MNIST for $E=1$ (all other settings were identical to those in the FC experiment). In each round, 7, 5, or 3 out of the 10 clients were randomly selected.
V-E Experiment 5: Imbalanced MNIST Dataset with the FC Model
@@
In this experiment, we assess the performance of $p$-FLARE in imbalanced scenarios of class distributions using the FC model. We partition the MNIST dataset into four distinct configurations: Distributing 1200 digit examples to 5 clients, with each client exclusively receiving examples corresponding to 2, 3, 4, or 5 different digit labels, while maintaining the test set unchanged. This setup introduces varying degrees of class imbalance, with each client possessing only 2 different digits (referred to as imbalance level 2), and dataset labels among clients do not overlap. The FC model is tested for $E=1$ and $R=0.001\%$. The Top-1 test accuracy results are presented in Fig.9. For the 2-level imbalance, after 1000 rounds, FedAvg achieves a test accuracy of 0.83, while $1$-FLARE achieves 0.74, and FFL and Error Correction achieve 0.67. For the 3-level imbalance, FedAvg achieves a test accuracy of 0.82, while $1$-FLARE achieves 0.7, and FFL and Error Correction achieve 0.54. In the case of 4-level imbalance, FedAvg achieves a test accuracy of 0.84, while $1$-FLARE achieves 0.74, and FFL and Error Correction achieve 0.62. Finally, for the 5-level imbalance, FedAvg achieves a test accuracy of 0.83, while $1$-FLARE achieves 0.73, and FFL and Error Correction achieve 0.59.
Figure 8: The evaluated loss on the test set for ”The Complete Works of William Shakespeare” experiment. FLARE is set with $R=0.001\%$, $p=\infty$, $\tau=0.05$, $c=1.01$. 10 clients are chosen, and we train for $E=1$ and batch size of 8. FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction, EF21, FedProx and FFL methods.
@@
V-F Experiment 6: Evaluation with reduced sparsity
Additional experiments are conducted to evaluate the performance of our proposed method at lower sparsity levels. Specifically, we included scenarios with $R=0.01\%$ and $R=0.1\%$ sparsity for the CNN experiments, as shown in Figures 10. Our findings indicate that as sparsity decreases, FLARE maintains superior performance compared to other error feedback algorithms. However, as sparsity increases, FLARE’s performance significantly surpasses the others, which is especially advantageous for dimensionality reduction applications. This trend is further validated by the theoretical performance bounds developed in Section IV, demonstrating that FLARE achieves a superior scaling rate with respect to sparsity.
(a) $E=4$
(b) $E=8$
(c) $E=16$
(d) $E=4$
(e) $E=8$
(f) $E=16$
@@ V-F Experiment 6: Evaluation with reduced sparsity
Figure 10: The test accuracy performance is compared for CNN model on digit MNIST with $R=0.01\%$ and $R=0.1\%$, $p=1$, $\tau=0.05$, $c=1.1$, 10 clients with $B=\infty$ and $E=4,8,16$. $p$-FLARE is compared with uncompressed FedAvg (benchmark for performance), Error Correction, EF21 and FedProx methods.
VI Conclusion